OpenStudy (anonymous):

which of the following are solutions to the equation below 8x^2-2x-2=1 a)x=3/4 or x=-1/2 B)x=7 or x=3 C) x=7/3 or x=3/2 D)x=7/3 pr x=-3/2 HELP PLEASE !!!!!!

OpenStudy (anonymous):

Well this looks simple. Just subsitute in the x values (the answer choices) and see if they equal 1.

OpenStudy (anonymous):

idk how to with the fractions. therefore im stuck

OpenStudy (anonymous):

To find the answer, just substitute in the answer choices. For example, if we were to substitute x=3/4 into the equation it would look like this: 8x^2-2x-2=1 8(3/4*3/4)-2(3/4)-2=1 8(9/16)-3/2-2=1 9/2-3/2-2=1 6/2-2=1 3-2=1 make sense?

OpenStudy (anonymous):

Maddie gave an..extended version ;)

OpenStudy (anonymous):

Thank you ! lol i get it now o: . i super suck at mathh -.- lmao